a solution containing 30% insecticide is to be mixed with solution containing 50% insecticide to make 200L of a solution containing 36% fungicide.how much of each solution should be used?
Answer by checkley71(8403)

.3X+.5(200-X)=.36*200
.3X+100-.5X=72
-.2X=72-100
-.2X=-28
X=-28/-.2
X=140L OF 30% FUNGICIDE IS NEEDED.
200-140=60L OF 50% FUNGICIDE IS NEEDED.
PROOF
.3*140+.5*60=72
42+30=72
72=72